OpenStudy (anonymous):

You drive a car 2.0 km to the gas station, 4.0 km to the shopping mall to let your mother off, and 4.5 km back home. The trip takes 20.0 min. What is your average speed? 3.2 m/s 4.38 m/s 8.75 m/s 0.525 m/s

OpenStudy (anonymous):

@Learner11

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@phi

OpenStudy (sophiaholder):

OKay so, S = distance/time

OpenStudy (sophiaholder):

So, you would want to add up how many kilometers you traveled first.

OpenStudy (anonymous):

that would be 10.5/20.00 right?

OpenStudy (sophiaholder):

Nope. If we are trying to find miles/seconds we need to convert them

OpenStudy (danjs):

<Speed> = ....

OpenStudy (anonymous):

oh

OpenStudy (danjs):

Do you remember calculating the 'average rate of change' given two points on a graph from algebra

OpenStudy (danjs):

slope of a secant line connecting the two points on the graph...

OpenStudy (sophiaholder):

Dan can probably explain this better than me, so I leave this one to him

OpenStudy (anonymous):

I have no Idea what your talking about my brain just died

OpenStudy (danjs):

It is the same as that, except it is a graph of (Total distance traveled) vs (total time traveled) for the y and x axis

OpenStudy (danjs):

To get the average speed - you take the initial value point (0 min, 0 km) and the final point (20 min , (4.5+2.0+4.0) km ) and calculate the slope of the line through those

OpenStudy (danjs):

In short, the average speed on an interval will be the ratio of the total distance traveled over the total time traveled

OpenStudy (anonymous):

@dan815

OpenStudy (danjs):

\[ave~speed = \frac{ distance ~traveled }{ time of traveling }\]

OpenStudy (danjs):

you will get km/min need to convert to m/s

OpenStudy (anonymous):

O.o

OpenStudy (dan815):

http://prntscr.com/8kh2ht

OpenStudy (dan815):

convert it into meters/second

OpenStudy (anonymous):

I DON'T UNDERSTAND x-x

OpenStudy (danjs):

\[\frac{ X~km }{ 1~\min }*\frac{ 10^3~m }{ 1~km }*\frac{ 1~\min }{ 60~s } = \frac{ X * 10^3 }{ 60 }*\frac{ m }{ s }\]

OpenStudy (anonymous):

omg

OpenStudy (dan815):

okay do you get that speed = d/t

OpenStudy (anonymous):

distance divided by time?

OpenStudy (dan815):

yes speed tells you the rate of distance change over time

OpenStudy (dan815):

you saw that 10.5 km traveled in 20 mins right

OpenStudy (dan815):

tell me if that part is clear or not, gotta know this to move on

OpenStudy (dan815):

10.5 = 2+4+4.5

OpenStudy (anonymous):

Everything